OpenStudy (anonymous):

Help required.... :( lim (1-sin theta)/ theta and theta approaches to 0

OpenStudy (anonymous):

\[\lim_{\theta \rightarrow 0}\frac{ 1-\sin \theta }{ \theta } \]

OpenStudy (kc_kennylau):

Hint: \(\lim\limits_{x\rightarrow0}\dfrac{\sin x}x=1\)

hartnn (hartnn):

in this case we need to consider both left hand limit and right hand limit know what those are ?

OpenStudy (mathmale):

According to the limit theorem for differences, Amtal's original limit can be broken into two separate limits: [lim as theta approaches zero of 1/theta] minus [lim as theta approaches zero of (sin theta)/theta]. The second limit should now be obvious. Hart is correct in that we need to consider both the left and right hand limits as theta approaches zero. Try it.

OpenStudy (anonymous):

So according to mathmale and hartnn guidance and suggestion, the two-side limits should be as following, you may try to draw y = 1/x to find out why \[\lim_{\theta \rightarrow 0^{+}} (\frac{ 1 }{ \theta } - 1) = +\infty\] and \[\lim_{\theta \rightarrow 0^{-}} (\frac{ 1 }{ \theta } - 1) = -\infty\]